• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 13:21:12 on December20,2003

    Entry number 114685

    keyword=Items Fixed In The Hall

    Last night there were problems with the left HRS dipole and with
    Q1. The dipole NMR probe would not lock and Q1 would only work
    for a subset of momentum.

    Today Scot and Ed came in and worked on the problems. Since we
    are using a limited momentum range, the NMR probe was set by hand
    for moments around 2 GeV. Scot tried to fix Q1, but the problem
    did not go away even after he changed all the boards, so we will
    need to avoid the problem moments during the run. (The problem
    acts like a stuck bit, so we have momentum gaps of 13 MeV.)

    Durning the access Ed recooled the shield. By doing this we should
    be able to run the magnet for the rest of the weekend without a recool.